OpenStudy (anonymous):

A rock weighs 130 N in air and has a volume of 0.00331 m3. What is its apparent weight when sub- merged in water? The acceleration of gravity is 9.8 m/s2 . Answer in units of N

OpenStudy (anonymous):

The buoyancy due to water offsets the rock's weight with a force equal to the product of displaced volume times density times gravitational acceleration, F = V rho g = (.00331m3)(1000kg/m3)(9.8N/kg)=30.4N, so the rock will weight this much less in water, or 130N-30.4N=99.6N. Note that g=9.8m/s2 is also g=9.8N/kg from force = mass x acceleration